# Task Assignment and Exploration Optimization for Low Altitude UAV Rescue via Generative AI Enhanced Multi-agent Reinforcement Learning

### v1

@@ 5.2.2 Area Selection for Exploration Optimization
The Hungarian algorithm is based on the bipartite graph matching theory. It continuously searches for augmenting trajectories and updates matching relationships to achieve the optimal matching solution. In this paper, the predetermined rescue area is divided into $\mathcal{B}$ subareas. UAVs will bear different costs when performing different tasks. The cost is as follows:
@@
Step 2 (Line 2-line 9). For each row (representing a UAV), the algorithm finds the optimal matching area, i.e., the optimal matching area for each UAV. If the optimal matching solution cannot be found directly, the algorithm adjusts the label value to ensure that a new matching relationship can be found. This process continues to iterate until all UAVs are assigned to the corresponding areas.
@@
1Initialize the label values and matching array: Set all the row labels $\alpha$ to the minimum value, i.e., $\alpha_{u}=\min\left(\mathbb{C}[u][1],...,\mathbb{C}[u][\mathcal{B}]\right)$, set all the column labels $\beta_{b}=0$.
2 for each UAV $u=1,2,...,U$ do
@@ 5.3.1 Motivation of Adopting GDM and MADDPG
The motivation for adopting GDM and MADDPG lies in their ability to enhance decision optimization in multi-agent environments, especially with limited offline training data, and GDM’s generative capabilities enable dynamic decision-making. Moreover, integrating GDM and MADDPG into UAV systems refines training processes and optimizes decision strategies to enhance coordination and adaptability in unknown environments. Furthermore, the Hungarian algorithm and GDM are innovatively integrated into MARL, which respectively realizes the computational complexity of the observation space and the generation capability of the action space, thus improving the observation and execution effects of multiple agents.
@@ 5.3.2 Markov Decision Process Modeling
Observations: In each time slot ${t}$, the UAV collects environmental observations $O_{i}=\{o_{i}^{1},o_{i}^{2},\cdots,o_{i}^{u}\}$. Then, the observation of an agent is defined as:
@@
5.3.3 UAVs Interacting with the Environment
In a multi-agent setup, the agent’s actor network generates specific actions based on its observations. These observations serve as inputs for the GDM’s inverse diffusion process, which incrementally predicts and denoises the sampled Gaussian noise to generate actions according to the current conditions. For agent ${u}$, the inverse denoising process aims to estimate the inverse distribution probability of a task, chosen from Gaussian noise $\mathbf{x}_{T}^{u}{\sim}N(0,I)$. If the probability $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is learned, it is feasible to get sampling $\mathbf{x}_{t}$ from a standard normal distribution and samples from $p(\mathbf{x}_{0})$ via the inverse denoising process. However, the estimation of $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is computationally complex in practice. Therefore, the objective is to approximate $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ using $\boldsymbol{p}_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{x}_{t},t))$. The probability from $\mathbf{x}_{T}$ to $\mathbf{x}_{0}$ can then be expressed as:
@@ 5.3.3 UAVs Interacting with the Environment
At each time step, the conditioning the model predicts the parameters of the Gaussian distribution, specifically the mean $\mu_{\theta}(\mathbf{x}_{t},t)$ and the covariance matrix $\boldsymbol{\Sigma}_{\theta}(\mathbf{x}_{t},t)$. Based on [47], incorporating conditional information $\mathbb{g}$ during the denoising process enables the model to be treated as a prediction model for noise, and the covariance matrix is fixed to $\boldsymbol{p}_{\theta}(\mathbf{x}_{0:T})$ and $\Sigma_{\theta}(\mathbf{x}_{t},\mathbb{g},t)=\beta_{t}{I}$. The mean is calculated as follows:
@@
Next, $\mathbf{x}_{T}$ sampled from $N(0,I)$, and through the inverse denoising process parameterized by $\theta$, we sample $\mathbf{x}_{t-1}\mid\mathbf{x}_{t}$ as follows:
@@
where $\boldsymbol{\epsilon}\sim N(0,I)$ represents a standard normal distribution, and $t=1,\cdots,T$.
@@
$\mathcal{L}_{t}=\mathbb{E}_{\mathbf{x}_{0},t,\boldsymbol{\epsilon}}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\sqrt{\overline{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\boldsymbol{\epsilon},t)\|^{2}\right].$ (20ak)
1The parameters of the critic-actor network and the inverse process of GDM
@@ 11                  An action selected by agent $u$ according to $o_{i}^{u}$
12                  Execute an action and obtain ${r}_{i}^{u}$ and the next $o_{i+1}^{u}$
@@ 17                         Update the memory
19
20                  for $u=0,1,\cdots,U$ do
21                         Samples are taken from the memory $(o_{i}^{u},A_{i}^{u},r_{i}^{u},o_{i+1}^{u},u),\forall{i}=1,2,\cdots,I$
23                        Update the parameters of the target network according to $\psi$
@@ 5.3.4 UAVs Interacting with Each Other
The experience information $(o_{i}^{u},A_{i}^{u},r_{i}^{u},o_{i+1}^{u},u)$ generated by the agent in the process of interacting with the environment is stored in the experience replay memory. Then, multiple agents share the experience by exchanging sample sets. Furthermore, the agent uses the time difference error to assign priorities to the experience information so that important data is sampled more frequently, thereby improving learning efficiency and performance. The critic network assesses the effectiveness of the action produced by the actor network through rewards, i.e., the impact of the generated action on the long-term reward. When each agent calculates the forward propagation of the critic network, it splices the observations of all agents, including itself, into the observation vector $O_{i}=\{o_{i}^{1},o_{i}^{2},...,o_{i}^{U}\}$, splices the actions of all agents into the action vector $A_{i}=\{A_{i}^{1},A_{i}^{2},...,A_{i}^{U}\}$, and uses $(O_{i},A_{i})$ as the input of the critic network and outputs a one-dimensional $Q$ value, i.e., $Q_{\theta_{Q}^{u}}(O_{i},A_{i})$. In other words, the agent uses the information of all other agents in the environment to centrally train its own evaluation network. Next, the target actor network calculates the action $A_{i+1}$ taken by the agent in the next observation through the sample $O_{i+1}$ in the experience replay memory, and then constructs the MSE loss function of $Q_{\theta_{Q}^{u}}(O_{i},A_{i})$ and $Q_{\theta_{Q}^{u}}(O_{i+1},A_{i+1})$ with the time difference error and uses gradient descent to update the parameter $\theta_{Q}^{u}$. The loss function and gradient formula are as follows:
$\begin{split}\min_{\theta_{Q}^{u}}\text{loss}&=\min_{\theta_{Q}^{u}}\mathbb{E}_{(o_{i}^{u},A_{i}^{u},r_{i}^{u},o_{i+1}^{u},u)\sim D}\Big{[}\Big{(}Q_{\theta_{Q}}(O_{i},A_{i})\\ &\quad\ -\big{(}r_{i}^{u}+\gamma Q_{\theta_{Q}}(O_{i+1},A_{i+1})\big{)}\Big{)}^{2}\Big{]}\end{split}$ (20al)
$\begin{split}\nabla_{\theta_{Q}^{u}}J(\theta_{Q}^{u})&=\nabla_{\theta_{Q}^{u}}\mathbb{E}_{(o_{i}^{u},A_{i}^{u},r_{i}^{u},o_{i+1}^{u},u)\sim D}\Big{[}\Big{(}Q_{\theta_{Q}}(O_{i},A_{i})\\ &\quad-\big{(}r_{i}^{u}+\gamma Q_{\theta_{Q}}(O_{i+1},A_{i+1})\big{)}\Big{)}^{2}\Big{]}\end{split}$ (20am)
@@ 5.3.5 Complexity Analysis of HG-MADDPG Algorithm
$\bullet$ Training stage: The computational complexity is given by $\mathcal{O}(2|\theta_{\mu}|+2|\theta_{Q}|+\mathbb{E}\delta|\theta_{\mu}|+\mathbb{E}\delta H\xi+U(2|\theta_{\mu^{\prime}}|+2|{\theta_{Q^{\prime}}}|))$, which can be broken down as follows [48]. The complexity of Parameters initialization of critic-actor network is $\mathcal{O}(2|\theta_{\mu}|+2|\theta_{Q}|)$, where $|\theta_{\mu}|$ and $|\theta_{Q}|$ are the number of parameters of the actor and critic networks, respectively. The complexity of Observation-action pair sampling is $\mathcal{O}(\mathbb{E}\delta|\theta_{\mu}|)$, where $\mathbb{E}$ is the total number of episodes. The complexity of Experience replay memory collection is $\mathcal{O}(\mathbb{E}\delta H\xi)$, where $H$ denotes the computational cost of the Hungarian algorithm, and $\xi$ denotes the complexity of interacting with the environment. The complexity of Target critic-actor network update is $\mathcal{O}(U(2|\theta_{\mu^{\prime}}|+2|{\theta_{Q^{\prime}}}|))$. The target network parameters are updated $U$ times. The space complexity denotes $\mathcal{O}(2|\theta_{\mu}|+2|\theta_{Q}|+\zeta(2|o_{i}^{u}|+|\mathrm{A}_{i}^{u}|+2))$, where $\zeta$ means the size of the memory. $|o_{i}^{u}|$ and $|\mathrm{A}_{i}^{u}|$ represent the dimension sizes of the observation and the action spaces, respectively. The complexity includes both the parameters of the neural network and the memory storing the tuples $(o_{i}^{u},A_{i}^{u},r_{i}^{u},o_{i+1}^{u},u)$.
@@ 6.1 Experiment Setup
  • MADDPG uses a framework of centralized training and distributed execution [27]. It exploit a centralized critic improves coordination by considering the actions of other agents.
  • MAPPO employs centralized training with decentralized execution [17]. It uses a clipping mechanism to stabilize learning and adjusts each agent’s policy independently.
@@ 6.2 Performance Evaluation
6.2.1 Convergence of training Procedure
@@
(a) Reward of different denoising steps.
(b) Reward of different batch sizes.
(c) Reward of different learning rates.
@@
(a) Queuing energy with weighting factor values.
(b) Queuing energy with different computing power.
(c) Queuing energy with different data sizes.
@@
(a) Latency under different computing powers.
(b) Latency under different numbers of GERs.
(c) Latency under different data sizes.
Figure 7: Stability of task completion latency under different data sizes, computing powers, and numbers of GERs.
(a) Latency under different computing powers.
(b) Latency under different numbers of GERs.
(c) Latency under different data sizes.
@@
(a) Area selection (colors) and two-dimensional trajectory.
(b) Three-dimensional trajectory.
Figure 9: Exploration optimization for UAVs.
@@ 6.2.3 Task Completion Latency
The experimental results shown in Fig. 8(a) and (b) illustrate how the task completion latency is affected by progressively increasing the number and computing power of GERs. This experiment was conducted with 3 UAVs and $V=0.5$. As the number and computing power of GERs increase, the task completion latency decreases. Furthermore, as shown in Fig. 8(c), while the computing power and number of GERs remain unchanged, as the task data size increases, the task queue becomes longer and the completion time increases accordingly. This shows that the above experimental results can clearly illustrate that the method proposed in this article is feasible and in line with common sense.
6.2.4 Altitude Exploration and Obstacle Avoiding
@@
Figure 10: Visualization of computing power.
6.2.5 Resource Visualization for Computing Power
Fig. 10 illustrates the distribution and utilization of computing power resources of GERs in a rescue area. The main purpose of this figure is to present the remaining computation resources of each GER. From a quantitative perspective, it can be observed that UAVs can autonomously select subareas and GERs based on their coordinates and available resources, as validated in Fig. 9. From a qualitative perspective, during rescue operations, emergency command decision-makers can quickly assess the available computation resources of GERs in the area through the visualization of their distribution and usage. This facilitates informed decision-making for the scheduling of rescue resources and the allocation of efforts.

### v2

@@ 5.2.2 Area Selection for Exploration Optimization
The exploration optimization involves trajectory generation and obstacle detection (see Section 3.5). Specifically, the UAV first selects an exploration area based on the Hungarian algorithm. Within the selected area, task offloading and GER selection are then performed using GDM-MADDPG. Once the GER is determined, the UAV’s trajectory is defined by its movement toward the selected GER to maximize the channel power gain (see Section 3.3). During this process, the UAV also conducts low-altitude obstacle detection to avoid potential collisions. The Hungarian algorithm is based on the bipartite graph matching theory. It continuously searches for augmenting trajectories and updates matching relationships to achieve the optimal matching solution. In this paper, the predetermined rescue area is divided into $\mathcal{B}$ subareas. UAVs incur varying costs when performing different tasks. The cost is as follows:
@@
Step 2 (Line 2-line 9). For each row (i.e., UAV) in the cost matrix, the algorithm finds the optimal matching area, i.e., the optimal matching area for each UAV. If the optimal matching solution cannot be found directly, the algorithm adjusts the label value to ensure that a new matching relationship can be found. This process continues to iterate until all UAVs are assigned to the corresponding areas.
@@
1Initialize the label values and matching array: Set all the row labels $\alpha$ to the minimum value, i.e., $\alpha_{u}=\min\left(\mathbb{C}[u][1],\ldots,\mathbb{C}[u][\mathcal{B}]\right)$, set all the column labels $\beta_{b}=0$.
2 for each UAV $u=1,2,\ldots,U$ do
@@ 5.3.1 Motivation of Adopting GDM and MADDPG
The motivation for adopting GDM and MADDPG lies in their ability to enhance decision optimization in multi-agent environments, especially with limited offline training data, and GDM’s generative capabilities enable dynamic decision-making [49]. Moreover, integrating GDM and MADDPG into UAV systems refines training processes and optimizes decision strategies to enhance coordination and adaptability in unknown environments. Furthermore, the Hungarian algorithm and GDM are innovatively integrated into MARL, which respectively realizes the computational complexity of the observation space and the generation capability of the action space, thus improving the observation and execution effects of multiple agents.
@@ 5.3.2 Markov Decision Process Modeling
Observations: In each time slot ${t}$, the UAV collects environmental observations $O_{i}=\{o_{i}^{1},o_{i}^{2},\ldots,o_{i}^{u}\}$. Then, the observation of an agent is defined as:
@@
5.3.3 UAVs Interacting with Environment
In a multi-agent setup, the agent’s actor network generates specific actions based on its observations. These observations serve as inputs for the GDM’s inverse diffusion process, which incrementally predicts and denoises the sampled Gaussian noise to generate actions according to the current conditions. For agent ${u}$, the purpose of the inverse diffusion process is to infer the task ratio $\mathbf{x}_{0}^{u}$ that is offloaded to the selected GER in the subarea from the Gaussian noise $\mathbf{x}_{T}^{u}{\sim}N(0,I)$. Since the number of GERs is random, the dimension of $\mathbf{x}_{0}^{u}$ depends on the maximum number of GERs in the subarea, that is, the dimension of the action space. If the probability $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is learned, it is feasible to get sampling $\mathbf{x}_{t}$ from a standard normal distribution and samples from $p(\mathbf{x}_{0})$ via the inverse denoising process. However, the estimation of $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is computationally complex in practice. Therefore, the objective is to approximate $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ using $\boldsymbol{p}_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{x}_{t},t))$. The probability from $\mathbf{x}_{T}$ to $\mathbf{x}_{0}$ can then be expressed as:
@@ 5.3.3 UAVs Interacting with Environment
The training of GDM involves optimizing the negative log-likelihood function of the training data. By adding conditional information $\mathbb{g}$, i.e., the agent’s observation $o_{i}^{u}$, during the denoising process. The airship assigns subareas to achieve load balancing of UAV exploration. Therefore, the UAV only needs to handle the offloading decision of the assigned subarea to achieve the minimum task completion delay and energy consumption, which can not only maintain the optimal decision of a single subarea and the global optimal allocation but also reduce the computational complexity. At each time step, the conditioning the model predicts the parameters of the Gaussian distribution, specifically the mean $\mu_{\theta}(\mathbf{x}_{t},t)$ and the covariance matrix $\boldsymbol{\Sigma}_{\theta}(\mathbf{x}_{t},t)$. Based on [50], incorporating conditional information $\mathbb{g}$ during the denoising process enables the model to be treated as a prediction model for noise, and the covariance matrix is fixed to $\boldsymbol{p}_{\theta}(\mathbf{x}_{0:T})$ and $\Sigma_{\theta}(\mathbf{x}_{t},\mathbb{g},t)=\beta_{t}{I}$. The mean is calculated as follows:
@@
Next, $\mathbf{x}_{T}$ sampled from $N(0,I)$. The sampling dimension depends on the number of GERs in each subarea. Then, through the inverse denoising process parameterized by $\theta$, we sample $\mathbf{x}_{t-1}\mid\mathbf{x}_{t}$ as follows:
@@
where $\boldsymbol{\epsilon}\sim N(0,I)$ represents a standard normal distribution, and $t=1,\ldots,T$.
@@
$\mathscr{L}_{t}=\mathbb{E}_{\mathbf{x}_{0},t,\boldsymbol{\epsilon}}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\sqrt{\overline{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\boldsymbol{\epsilon},t)\|^{2}\right].$ (20ak)
1The parameters of the actor-critic network and the inverse process of GDM
@@ 11                  An action selected by agent $u$ according to $o_{i}^{u}$
12             end for
13
14            Execute action and obtain ${r}_{i}^{u}$ and the next $o_{i+1}^{u}$
@@ 19                   Update the memory
20                  for $u=1,2,\ldots,U$ do
21                         Samples are taken from the memory $(o_{i}^{u},A_{i}^{u},r_{i}^{u},o_{i+1}^{u},u),\forall{i}=1,2,\ldots,I$
23                   end for
24                  Update the parameters of the target network according to $\psi$
@@
26
@@ 5.3.4 UAVs Interacting with Each Other
The experience information $(o_{i}^{u},A_{i}^{u},r_{i}^{u},o_{i+1}^{u},u)$ generated by the agent in the process of interacting with the environment is stored in the experience replay memory. Then, multiple agents share the experience by exchanging sample sets. Furthermore, the agent uses the time difference error to assign priorities to the experience information so that important data is sampled more frequently, thereby improving learning efficiency and performance. The critic network assesses the effectiveness of the action produced by the actor network through rewards, i.e., the impact of the generated action on the long-term reward. When each agent calculates the forward propagation of the critic network, it splices the observations of all agents, including itself, into the observation vector $O_{i}=\{o_{i}^{1},o_{i}^{2},\ldots,o_{i}^{U}\}$, splices the actions of all agents into the action vector $A_{i}=\{A_{i}^{1},A_{i}^{2},\ldots,A_{i}^{U}\}$, and uses $(O_{i},A_{i})$ as the input of the critic network and outputs a one-dimensional $Q$ value, i.e., $Q_{\theta_{Q}^{u}}(O_{i},A_{i})$. In other words, the agent uses the information of all other agents in the environment to centrally train its own evaluation network. Next, the target actor network calculates the action $A_{i+1}$ taken by the agent in the next observation through the sample $O_{i+1}$ in the experience replay memory, and then constructs the MSE loss function of $Q_{\theta_{Q}^{u}}(O_{i},A_{i})$ and $Q_{\theta_{Q}^{u}}(O_{i+1},A_{i+1})$ with the time difference error and uses gradient descent to update the parameter $\theta_{Q}^{u}$. The loss function and gradient formula are as follows:
$\begin{split}\min_{\theta_{Q}^{u}}\text{loss}&=\min_{\theta_{Q}^{u}}\mathbb{E}_{(o_{i}^{u},A_{i}^{u},r_{i}^{u},o_{i+1}^{u},u)\sim D}\Big{[}\Big{(}Q_{\theta_{Q}}(O_{i},A_{i})\\ &\quad\ -\big{(}r_{i}^{u}+\gamma Q_{\theta_{Q}}(O_{i+1},A_{i+1})\big{)}\Big{)}^{2}\Big{]},\end{split}$ (20al)
$\begin{split}\nabla_{\theta_{Q}^{u}}J(\theta_{Q}^{u})&=\nabla_{\theta_{Q}^{u}}\mathbb{E}_{(o_{i}^{u},A_{i}^{u},r_{i}^{u},o_{i+1}^{u},u)\sim D}\Big{[}\Big{(}Q_{\theta_{Q}}(O_{i},A_{i})\\ &\quad-\big{(}r_{i}^{u}+\gamma Q_{\theta_{Q}}(O_{i+1},A_{i+1})\big{)}\Big{)}^{2}\Big{]},\end{split}$ (20am)
@@ 5.3.5 Complexity Analysis of HG-MADDPG Algorithm
$\bullet$ Training stage: The computational complexity is given by $\mathcal{O}(2|\theta_{\mu}|+2|\theta_{Q}|+\mathbb{E}\delta|\theta_{\mu}|+\mathbb{E}\delta H\xi+U(2|\theta_{\mu^{\prime}}|+2|{\theta_{Q^{\prime}}}|))$, which can be broken down as follows [51]. The complexity of Parameters initialization of actor-critic network is $\mathcal{O}(2|\theta_{\mu}|+2|\theta_{Q}|)$, where $|\theta_{\mu}|$ and $|\theta_{Q}|$ are the number of parameters of the actor-critic networks, respectively. The complexity of Observation-action pair sampling is $\mathcal{O}(\mathbb{E}\delta|\theta_{\mu}|)$, where $\mathbb{E}$ is the total number of episodes. The complexity of Experience replay memory collection is $\mathcal{O}(\mathbb{E}\delta H\xi)$, where $H$ denotes the computational cost of the Hungarian algorithm, and $\xi$ denotes the complexity of interacting with the environment. The complexity of Target actor-critic network update is $\mathcal{O}(U(2|\theta_{\mu^{\prime}}|+2|{\theta_{Q^{\prime}}}|))$. The target network parameters are updated $U$ times. The space complexity denotes $\mathcal{O}(2|\theta_{\mu}|+2|\theta_{Q}|+\zeta(2|o_{i}^{u}|+|\mathrm{A}_{i}^{u}|+2))$, where $\zeta$ means the size of the memory. $|o_{i}^{u}|$ and $|\mathrm{A}_{i}^{u}|$ represent the dimension sizes of the observation and the action spaces, respectively. The complexity includes both the parameters of the neural network and the memory storing the tuples $(o_{i}^{u},A_{i}^{u},r_{i}^{u},o_{i+1}^{u},u)$.
@@ 6.1 Experiment Setup
  • MADDPG is a decentralized actor–centralized critic algorithm designed primarily for environments with continuous action spaces [29]. The algorithm enables each agent to learn its own deterministic policy while leveraging centralized training to learn global information, which facilitates coordination among agents in partially observable settings. However, MADDPG can face challenges in training stability and scalability as the number of agents grows.
  • MAPPO is a policy-gradient-based method adapted from PPO, which uses clipped surrogate objectives to improve training stability and sample efficiency [19]. MAPPO supports centralized training with decentralized execution and works well in both discrete and continuous action spaces.
@@ 6.2 Performance Evaluation
6.2.1 Convergence of training Process
@@
(a) Reward vs. denoising step.
(b) Reward vs. batch size.
(c) Reward vs. learning rate.
@@
(a) Queuing energy vs. weighting factor.
(b) Queuing energy vs. computing power.
(c) Queuing energy vs. data size.
@@
(a) Latency vs. computing power.
(b) Latency vs. number of GERs.
(c) Latency vs. data size.
Figure 7: Stability of task completion latency under different data sizes, computing powers, and number of GERs.
(a) Latency vs. computing power.
(b) Latency vs. number of GERs.
(c) Latency vs. data size.
@@ 6.2.3 Task Completion Latency
[…]
6.2.4 Altitude Exploration and Resource Visualization
@@
(a) Subarea (colors) and 2D trajectory.
(b) 3D trajectory and obstacle avoidance.
(c) Visualization of computing power.
Figure 9: Exploration optimization and GER distribution.
Fig. 9(c) illustrates the distribution and utilization of computing power resources of GERs in a rescue area. The main purpose of this figure is to present the remaining computation resources of each GER. From a quantitative perspective, it can be observed that UAVs can autonomously select subareas and GERs based on their coordinates and available resources, as validated in Fig. 9(a). From a qualitative perspective, during rescue operations, emergency command decision-makers can quickly assess the available computation resources of GERs in the area through the visualization of their distribution and usage. This facilitates informed decision-making for the scheduling of rescue resources and the allocation of efforts.